### v1

# Agentic Search in the Wild: Intents and Trajectory Dynamics from 14M+ Real Search Requests

Abstract.
LLM-powered search agents are increasingly being used for multi-step information seeking tasks, yet the IR community lacks empirical understanding of how agentic search sessions unfold and how retrieved evidence is used. This paper presents a large-scale log analysis of agentic search based on 14.44M search requests (3.97M sessions) collected from DeepResearchGym, i.e. an open-source search API accessed by external agentic clients. We sessionize the logs, assign session-level intents and step-wise query-reformulation labels using LLM-based annotation, and propose Context-driven Term Adoption Rate (CTAR) to quantify whether newly introduced query terms are traceable to previously retrieved evidence. Our analyses reveal distinctive behavioral patterns. First, over 90% of multi-turn sessions contain at most ten steps, and 89% of inter-step intervals fall under one minute. Second, behavior varies by intent. Fact-seeking sessions exhibit high repetition that increases over time, while sessions requiring reasoning sustain broader exploration. Third, agents reuse evidence across steps. On average, 54% of newly introduced query terms appear in the accumulated evidence context, with contributions from earlier steps beyond the most recent retrieval. The findings suggest that agentic search may benefit from repetition-aware early stopping, intent-adaptive retrieval budgets, and explicit cross-step context tracking. We plan to release the anonymized logs to support future research.
Agentic Search, Query Log Analysis, Deep Research, Search Intent
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/26/07conference: The 49th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 20–24, 2026; Melbourne — Naarm, Australia
1. Introduction
Information retrieval is shifting from human-initiated search into agentic search (Asai et al., 2024; Nakano et al., 2022; Schick et al., 2023; Yao et al., 2023), where LLM-powered agents plan and execute multi-step information seeking with retrieval tools. Instead of issuing a single query and consuming a ranked list, an agent iteratively reformulates queries, retrieves evidence, and updates subsequent queries based on retrieved context. While agent capabilities are increasingly demonstrated on controlled benchmarks (Jin et al., 2025; Mialon et al., 2023; Wu et al., 2025), the benchmark scores alone do not reveal how agents’ queries evolve across steps, or how retrieved context shapes subsequent queries.
The aforementioned questions matter for practical system design. Agents may waste retrieval budget through repetitive or overly narrow reformulations, fail to explore alternative facets, or under-use evidence accumulated across steps. Understanding session structure can inform query-policy control, and quantifying evidence reuse can guide budget allocation and evaluation design. As agents consume results programmatically, leaving no observable trace of what they found useful, the logs lack implicit feedback signals (e.g., direct clicks) that anchor traditional behavioral inference. This creates a measurement gap. We can observe sequences of queries and retrieved evidence, but it remains unclear how sessions unfold, how behavior differs by intent, what reformulation moves dominate, and whether agents incorporate evidence from earlier steps.
Figure 1. Intent–trajectory structure of agentic search logs.
To address this gap, we analyze agentic search at two complementary levels: what the agent is trying to accomplish in a session (session-level intent), and how the agent pursues that goal through step-wise search actions (trajectory-level query reformulation). Figure 1 illustrates this structure on a short example session. To operationalize these levels, we develop a measurement framework with three components: (1) LLM-based annotation pipelines that assign interpretable intent and trajectory labels to sessions and step-pairs following standard taxonomies; (2) offline replay of logged queries to reconstruct the retrieved evidence available at each step; and (3) Context-driven Term Adoption Rate (CTAR), a metric we introduce to quantify whether newly introduced query terms can be lexically traced to retrieved evidence, including contributions from steps beyond the most recent retrieval.
We apply this framework to logs collected from DeepResearchGym (DRGym) (Coelho et al., 2025), i.e. a research-oriented reproducible search API accessed by external agentic clients. With permission from the DRGym organizers, we study 14.44M logged search requests spanning six months, which we sessionize into 3.97M sessions. This provides an at-scale view of autonomous agents operating in the wild under a shared retrieval backend.
Our results answer the practical questions raised above about budget waste, reformulation behavior, and evidence reuse in multi-step agentic search. We find that over 90% of multi-turn sessions contain at most ten steps, and 89% of inter-step intervals fall under one minute. Retrieval depth (i.e. the number of documents requested per query) is largely static, suggesting agents treat it as a fixed parameter rather than adapting within sessions. Behavior also varies by intent. Fact-seeking sessions exhibit the highest repetition, which increases over time, indicating that agents can enter near-duplicate loops when retrieval is unproductive, while sessions requiring reasoning sustain broader exploration throughout. We also find that agents frequently incorporate terms from previously retrieved evidence, with measurable contributions from earlier steps beyond the most recent retrieval.
Our contributions can be summarized as follows:
  • We provide a large-scale behavioral characterization of agentic search from a reproducible search infrastructure (14.44M search requests, 3.97M sessions), offering an at-scale view of autonomous agents operating in the wild.
  • We introduce CTAR, a metric for quantifying evidence-conditioned query evolution, and use it to demonstrate cross-step evidence reuse beyond the most recent retrieval.
  • We provide practical design takeaways from the logs, including repetition-aware stopping, intent-adaptive retrieval budgeting, and cross-step context tracking.
We plan to release an anonymized version of the logs to support reproducibility (details in Section 3).
2. Related Work
Human Search Behavior and Log Analysis:
Large-scale query logs have long been used to study search behavior in the wild (Dumais et al., 2014; Jansen et al., 1998; Silverstein et al., 1999), offering scalable, behavior-grounded signals for characterizing session dynamics and query reformulation beyond what offline benchmarks capture. A core theme is within-session learning. Eickhoff et al. (Eickhoff et al., 2014) trace how newly introduced terms relate to evidence observed before reformulation (e.g., SERP snippets and visited pages), alongside complementary work on interpreting implicit feedback such as clicks and dwell time (Agichtein et al., 2006; Fox et al., 2005; Joachims et al., 2005). Exploratory search and navigation studies further document differences in branching and interaction patterns across users and information needs (Marchionini, 2006; Teevan et al., 2004; White and Drucker, 2007), while sessionization analyses examine how sessions begin and end in practice (Jones and Klinkner, 2008). We adopt this evidence-traceability perspective for autonomous agents and operationalize it using retrieved evidence text, enabling systematic comparisons of agent behaviors across intents and guiding design choices such as retrieval budgeting and cross-step context management. While a small line of work compares humans and agents directly (Wang et al., 2025a, b; Zhou et al., 2025), these comparisons are often task-specific or simulation-based, motivating complementary large-scale log analyses of how autonomous agents search across sessions in the wild.
LLM Interaction Platforms and Usage Logs:
Recent efforts analyze large-scale interaction data from LLM systems and evaluation platforms. Chatbot Arena (LMSYS LLM Arena) aggregates pairwise preference votes (Chiang et al., 2024), and LMSYS-Chat-1M releases one million multi-model conversations collected in the wild (Zheng et al., 2024). OpenAI reports how people use ChatGPT at scale (OpenAI, 2025b), and Anthropic presents privacy-preserving analyses of millions of Claude conversations to characterize economic task usage (Handa et al., 2025). SciArena extends the Arena protocol to scientific literature-grounded tasks and provides a corresponding benchmark (Zhao et al., 2025). These works capture usage and preference signals, but typically do not expose tool-level retrieval traces (queries, evidence, step-wise search decisions) needed to study agentic search behavior and within-session evidence reuse.
(a) Geographic distribution of requests.
(b) Pairwise query cos similarity (100k sample).
(c) Query frequency distribution (log–log).
Figure 2. Representativeness and diversity of the DRGym logs.
Agentic Search Modeling, Benchmarks, and Infrastructures:
Recent systems enabling LLMs to plan multi-step interactions with retrieval tools have shifted IR toward agentic workflows (Asai et al., 2024; Nakano et al., 2022; Schick et al., 2023; Yao et al., 2023). Benchmarks for tool-using agents include WebShop (Yao et al., 2022), WebArena (Zhou et al., 2023), AgentBench (Liu et al., 2023), and large-scale tool-use evaluation such as ToolLLM/ToolBench (Qin et al., 2023). DeepResearchGym (DRGym) provides an open-source sandbox with a reproducible search API and evaluation protocol for deep research systems (Coelho et al., 2025). Early analyses have begun to formalize agent behaviors. Jin et al. (Jin et al., 2025) link beneficial reasoning patterns to gains on GAIA (Mialon et al., 2023) and WebWalker (Wu et al., 2025); complementary efforts propose taxonomies and risk frameworks, such as ST-WebAgentBench (Levy et al., 2025) and the Agentic AI Security Scoping Matrix (Brown and Saner, 2025). However, benchmark scores alone provide limited visibility into how agents search in practice, and prior human–agent comparisons suggest differences in query breadth and context use that are difficult to diagnose without session-level traces (Wang et al., 2025a, b; Zhou et al., 2025). Most prior work focuses on benchmarking and system design, whereas we measure behavior at scale from real logs, and quantify evidence traceability via CTAR.
3. Data and Log Processing
In this section, we start by providing an overview of DRGym to better contextualize our analysis. Then, we describe the query log we have been given access to, following with the preprocessing and session segmentation (sessionization) pipeline used to convert raw requests into sessions.
3.1. DRGym Log Overview
DRGym serves requests from external agentic clients, capturing diverse usage patterns and interaction styles. The API is model-agnostic; although logs do not record the underlying agent implementation, requests originate from independent client systems operating under the same retrieval infrastructure.
The backend performs dense retrieval (Jayaram Subramanya et al., 2019; Karpukhin et al., 2020) over two large-scale English corpora, ClueWeb22-A-EN (Overwijk et al., 2022) and FineWeb (Penedo et al., 2024). The DRGym paper describes a retrieval API with a /search endpoint for ranked retrieval (Coelho et al., 2025). Operating over static web snapshots (instead of a changing live index) enables re-issuing queries under fixed corpora for consistent retrieval behavior across experiments.
Consistent with this design, each log entry records the timestamped query and request parameters (Table 1), including retrieval depth and an ANN search-budget parameter (Jayaram Subramanya et al., 2019). For privacy, IP addresses are anonymized and used only for coarse client-level aggregation (grouping, sessionization, and country-level reporting), never for user identification or fine-grained geolocation.
Scale and Coverage:
The logs span 2025-06 to 2025-12 and contain 14.44 million requests. After preprocessing and sessionization (Section 3.2), we obtain 3.97M sessions. Requests originate from 558 anonymized client IPs across 25 countries. Figure 2(a) shows the geographic coverage, with the largest shares from China and the United States, followed by Hong Kong and Iceland. Overall traffic is substantial, peaking at 2.49 million requests in a single week.
Table 1. Fields recorded in the search_logs table.
Field Description
id Auto-incremented unique identifier for each request.
ip_address Client IP address (anonymized for analysis).
query_text Query string submitted to the API.
num_of_docs Number of documents requested for retrieval.
complexity ANN retrieval complexity controlling search budget.
dataset Corpus used for retrieval (ClueWeb22 or FineWeb).
timestamp Timestamp of the logged request.
Semantic Diversity:
[…]
Query-Level Repetition:
Figure 2(c) shows a long-tailed query frequency distribution. Most distinct queries are quite rare, while only a small set of queries repeats often. In particular, 53.89% of distinct queries occur at most three times, including 38.38% singleton queries, and the top-10 and top-100 most frequent queries account for only 0.59% and 1.51% of all requests, respectively.
Taken together, the broad geographic coverage, low average semantic similarity, and long-tailed frequency spectrum suggest that the DRGym stream reflects a realistic and diverse mix of information needs, rather than a narrow set of repeatedly executed prompts. To further validate that this diversity is not an artifact of surface-level paraphrasing of popular benchmarks, we next quantify the overlap between logged queries and public benchmarks.
We measure semantic overlap between log queries and four commonly used agentic benchmarks: GAIA (Mialon et al., 2023), FRAMES (Krishna et al., 2025), HLE (Phan et al., 2025), and WebWalkerQA (Wu et al., 2025). We consider a sample of 1 million queries from the logs, and encode all benchmark queries using Qwen3-Embedding-0.6B. Then, we count log queries exceeding a cosine similarity threshold of 0.7 to any benchmark query. As shown in Table 2, benchmark-similar queries constitute less than 0.4% of the sample across all four benchmarks combined. Overlap is lowest for GAIA, and highest for WebWalker, whose web traversal queries more closely resemble natural search formulations. Overall, these results suggest the logs reflect diverse open-ended usage rather than concentrated benchmark execution.
Table 2. Semantic overlap between agentic benchmarks and a 1M log query sample (cosine similarity $\geq 0.7$).
Benchmark Bench. Queries Hits in Sample % of Sample
GAIA 103 27 0.00%
FRAMES 824 879 0.09%
HLE 2,158 616 0.06%
WebWalkerQA 680 2,219 0.22%
Total 3,765 3,741 0.37%
To support reproducibility and enable further research, we will release the cleaned and anonymized logs associated with this study. Prior to release, we will remove direct identifiers (e.g., IP addresses) and apply standard PII scrubbing on free-text fields, releasing only the fields needed to reproduce our analyses with anonymized session IDs. This release has already received formal approval from the DRGym organizers. We will document the anonymization procedure and residual risks in the dataset card, following prior large-scale LLM interaction log releases and privacy-preserving analyses (Handa et al., 2025; Zheng et al., 2024). After the initial publication, the dataset will be updated as additional logs are collected and validated.
3.2. Log Preprocessing and Sessionization
We first remove malformed entries (e.g., empty queries), internal testing traffic, and outlier repetition bursts, before segmenting the remaining stream into sessions.
Although standard sessionization often relies on fixed time-gap heuristics, agentic requests can arrive in fast parallel patterns (Nie et al., 2025), making a pure temporal cutoff unreliable. We therefore sessionize with a semantic-continuity criterion combined with an explicit temporal constraint. Concretely, for each IP we maintain active sessions and assign an incoming query to the most semantically continuous active session when the continuity score exceeds a threshold; otherwise we start a new session. We additionally impose a 10-minute hard cutoff between consecutive queries within a session, reflecting faster interaction loops than the conventional 30-minute rule for human logs (Jones and Klinkner, 2008; Silverstein et al., 1999). Among classifier-predicted continuous pairs, only 0.92% have gaps exceeding 10 minutes.
The aforementioned pipeline yields 3.97M sessions. Manual spot-checks confirm that the resulting sessions are generally coherent. Full procedural details (i.e., continuity model, thresholds, and validation) are provided in Appendix A.
4. Methodology
To address the questions motivating this work, we require measurements at two levels, which are session-level intent (what type of information need drives the session) and trajectory-level reformulation (how queries change from step to step). We also require a way to assess whether retrieved evidence influences subsequent queries. For intent and trajectory labeling, we use standard LLM-as-a-judge pipelines (Li et al., 2024; Zheng et al., 2023). For evidence influence, we introduce a new metric, namely the Context-driven Term Adoption Rate (CTAR).
We segment the log into sessions $\mathcal{S}$, where each session $s=(q_{1},\ldots,q_{|s|})$ is an ordered sequence of timestamped queries. Retrieval depth is denoted by $K$, corresponding to the logged parameter num_of_docs. We analyze behavior at three granularities: global (corpus-wide), session-level (intent-conditioned), and trajectory-level (adjacent query pairs within a session $q_{k}\rightarrow q_{k+1}$).
4.1. LLM-based Intent and Trajectory Labeling
Session-Level Intent:
Different information needs may induce different search strategies. For instance, a user seeking a factual answer may behave differently from one debugging a procedure or reasoning through a complex question. To test whether agentic search exhibits such intent-conditioned structure, we label each session with an intent category. We adopt a three-way taxonomy from web search goal modeling (Broder, 2002; Eickhoff et al., 2014; Rose and Levinson, 2004): Declarative (fact retrieval), Procedural (method execution), and Reasoning (complex synthesis). Since $q_{1}$ is often already a reformulation, we assign intent from the whole session text.
Trajectory-Level Reformulation:
Intent alone does not reveal how agents iterate within a session. An agent might narrow its query, broaden it, pivot to a related facet, or retry with a near-identical phrasing. These reformulation patterns have implications for retrieval efficiency: excessive repetition wastes budget, while a lack of exploration may leave relevant facets unexamined. To capture these dynamics, we label each adjacent query pair ($q_{k}\rightarrow q_{k+1}$) with a trajectory type grounded in prior reformulation taxonomies (Boldi et al., 2011; Huang and Efthimiadis, 2009): Specialization (narrowing by adding constraints), Generalization (broadening by relaxing constraints), Exploration (within-topic facet pivots) and Repetition (identical or near-duplicate reformulations). Representative examples are provided in Appendix C.
Implementation:
We implement labeling with gpt-5-nano (OpenAI, 2025a). We annotate multi-turn sessions with $|s|\!\in\![2,10]$ for intent (one label per session) and all adjacent pairs for trajectories (one label per pair). We focus on this range because our sessionization analysis (Section 5) reveals that it covers 90.32% of all multi-turn traffic, representing the core behavior of current agents. To assess labeling robustness, we compare labels from two models (gpt-5-nano and gemini-3-flash-preview (Google, 2025)) on a 2000-pair random subset, achieving 95.15% agreement; the remaining disagreements are spread across categories rather than concentrated in any single label. Prompts are provided in Appendix D. Unless otherwise noted, analyses from Section 6 onward use a random subset of labeled multi-turn sessions under the annotation budget (excluding single-query sessions and outlier long-tail sessions, as described in Section 5). We also compute auxiliary metrics used throughout the paper, each defined at first use with a summary in Appendix B.
Table 3. Session-level descriptive statistics by intent type. Formulas for less standard metrics are in Appendix B.
Statistic Metric Declarative (fact-seeking) Procedural (how-to) Reasoning (analytical)
(Sample N / Ratio) (99.8k / 88.64%) (4.5k / 3.96%) (8.3k / 7.41%)
Mean Session Length 4.03 3.81 4.03
Retrieval Depth ($K$) 7.70 37.34 24.99
Query Length (whitespace terms) 7.59 10.58 12.69
Initial-Final Gap 0.21 0.22 0.28
Median Total Duration (s) 40.00 26.00 31.00
Step Latency (s) 17.00 13.00 14.00
4.2. Context-driven Term Adoption Rate (CTAR)
Agentic search iterates between retrieval and query formulation. Evidence returned at step $k$ may shape how the agent revises the next query. Yet agentic logs provide no direct signal of what the agent actually attended to in retrieved documents, making evidence use hard to observe. We therefore ask a more tractable traceability question: when the agent introduces new query terms at step k+1k{+}1, do those terms appear in the evidence it has already seen? This aligns with evidence-traceability perspectives in human log analysis (Eickhoff et al., 2014) and searching-as-learning research (Eickhoff et al., 2015; Rieh et al., 2016; Urgo and Arguello, 2022), but has not been systematically studied for autonomous agents.
We formulate this idea through Context-driven Term Adoption Rate (CTAR), i.e. the fraction of newly introduced query terms that can be lexically traced to retrieved evidence. We use exact-match tracing rather than semantic similarity because it is interpretable without threshold tuning, robust across domains and query styles, and conservative, i.e., semantic variants would typically yield higher rates by crediting paraphrases and near matches.
Let $Terms(x)$ denote the set of unique, lowercased, and non-stopword tokens that can be extracted from text $x$. For a trajectory $(q_{k}\rightarrow q_{k+1})$, the set of newly introduced terms is:
(1) $NewTerms(q_{k+1},q_{k})=Terms(q_{k+1})\setminus Terms(q_{k}).$
Let $E_{k}$ denote the textual evidence available at step $k$. Since raw logs do not store retrieved documents, we reconstruct $E_{k}$ by querying the DRGym API (which guarantees reproducibility) using the original logged parameters. We consider two context definitions:
(2) $C_{k}^{\text{last}}$ $=Terms(E_{k}),$
(3) $C_{k}^{\text{agg}}$ $=\bigcup_{i=1}^{k}Terms(E_{i}).$
CTAR is the fraction of new terms appearing in the chosen context:
(4) $\text{CTAR}^{(\cdot)}_{k}=\frac{\left|NewTerms(q_{k+1},q_{k})\cap C_{k}^{(\cdot)}\right|}{\left|NewTerms(q_{k+1},q_{k})\right|},\quad(\cdot)\in\{\text{last},\text{agg}\}.$
In the previous equation, if $NewTerms(q_{k+1},q_{k})=\emptyset$ or $C_{k}^{(\cdot)}=\emptyset$, we set $\text{CTAR}^{(\cdot)}_{k}=0$. In summary, CTAR quantifies the degree to which query evolution is lexically grounded in retrieved evidence. $\text{CTAR}^{last}$ captures adoption from the immediately preceding step, while $\text{CTAR}^{agg}$ captures adoption from any prior step, enabling us to test whether agents integrate evidence across the full session or rely only on the most recent retrieval. Comparing the two allows quantifying the contribution of earlier context, and identifying session types or phases where evidence integration breaks down.
5. Aggregate Session Statistics
Session Length and Structural Composition:
The corpus comprises 3.97M unique search sessions with a skewed length distribution (Figure 3, left). Nearly half (47.77%) are single-query sessions (likely one-off API calls or simple lookups), with manual inspection suggesting that these are predominantly Declarative or Procedural queries. Among multi-turn sessions, 90% have length $\leq 10$, indicating that most multi-turn sessions in the log resolve within a few iterations, though some extend considerably further.
For reference, human web search logs are slightly shorter on average (1.7 queries per session) and include a substantially larger fraction of single-query sessions (77.6%) (Eickhoff et al., 2014; Silverstein et al., 1999), suggesting that agentic systems engage in more extended information-seeking episodes. We focus subsequent analyses on multi-turn sessions.
Figure 3. Left: distribution of session length (number of queries per session). Right: distribution of step-wise time intervals between consecutive requests.
Temporal Dynamics and Interaction Speed:
Within-session intervals are short for most steps, where 56.12% fall within 0–10 seconds, and 89.21% are under one minute (Figure 3, right). Intervals are heavy-tailed, reflecting occasional long-latency steps due to system or pipeline delays.
For reference, prior human log studies report median dwell times of several minutes for knowledge-acquisition intents (Eickhoff et al., 2014). While dwell time and session duration are not directly comparable to our inter-step intervals, the contrast highlights the faster iterative pacing typical of agentic search.
Retrieval Depth and Parameter Stability:
Retrieval depth is concentrated at fixed values $K\in\{1,5,10\}$, with only 8.36% of sessions using other values. Furthermore, only 1.35% of sessions vary $K$ across steps. Since DRGym supports $1\leq K\leq 100$, this suggests that many agents treat retrieval count as effectively hard-coded rather than adapted within a session.
Table 4. Descriptive statistics by trajectory type. Formulas for less standard metrics are in Appendix B.
Statistic Metric Specialization Generalization Exploration Repetition
(Sample N / Ratio) (73.8k / 21.76%) (30.6k / 9.02%) (125.8k / 37.07%) (109.1k / 32.15%)
Narrowing Broadening Facet pivots Near-duplicates
Mean Dense Similarity 78.18% 80.07% 55.08% 96.70%
Jaccard Similarity 40.13% 44.66% 25.10% 81.63%
Result Overlap 22.58% 23.47% 7.35% 78.23%
Delta Query Length ++1.47 -2.42 ++0.04 -0.08
Median Step Latency (s) 7.00 8.00 14.00 6.00
6. Intent-Conditioned Session Behavior
Using the LLM-as-a-judge pipeline described in Section 4, we label each multi-turn session as either being Declarative (fact-seeking), Procedural (how-to or step-by-step tasks), or Reasoning (comparative, analytical or multi-hop questions). Declarative dominates (88.64%), followed by Reasoning (7.41%) and Procedural (3.96%). Table 3 summarizes the session-level behavior, reporting medians for time-based measures due to heavy tails and means for count and semantic measures.
Beyond standard length and timing, we characterize sessions with two additional measures. Retrieval Depth summarizes per-step $K$ at the session level, and Initial-Final Gap measures semantic drift from first to last query $1-\cos(q_{1},q_{|s|})$ using Qwen3-Embedding-0.6B (Zhang et al., 2025). Formal definitions are in Appendix B.
Declarative sessions use the shallowest retrieval yet incur the highest interaction costs. This suggests that limited evidence per step forces agents into more iterations to locate and verify information. The pattern contrasts with human fact-finding, where users issue fewer and shorter queries (Dumais et al., 2014; Eickhoff et al., 2014; Silverstein et al., 1999). Agents instead phrase queries as full constraint-bearing questions, consistent with iterative verification behavior (Jin et al., 2025).
Procedural sessions show the opposite pattern. Deeper retrieval accompanies a more semantically stable progression, suggesting that broader evidence coverage reduces the need for iterative refinement. Queries within these sessions are longer than Declarative ones, consistent with prior studies of procedural search reporting similar characteristics (Eickhoff et al., 2014).
Finally, we observe that Reasoning sessions match Declarative in turn count but differ in how queries evolve. They show the largest semantic drift and longest queries, while retrieval depth is moderate. The distinguishing signal for Reasoning lies in within-session query reformulation rather than session duration or retrieval depth.
7. Trajectory Moves and Topologies
In this section, we analyze how agents revise queries step by step, a defining characteristic of agentic search that exposes intermediate decision-making and supports more fine-grained analysis than single-shot querying.
7.1. Trajectory Types and Properties
Table 5. Distribution of trajectory types across all step-wise transitions within sessions of each intent.
Trajectory Session Intent
Declarative Procedural Reasoning
Specialization 21.12% 27.99% 26.35%
Generalization 9.07% 10.12% 7.89%
Exploration 36.12% 39.04% 47.57%
Repetition 33.69% 22.85% 18.19%
We follow the labeling procedure in Section 4.1, classifying each adjacent pair $(q_{k}\rightarrow q_{k+1})$ as Specialization (narrowing by adding constraints), Generalization (broadening by relaxing constraints), Exploration (within-topic facet pivots), or Repetition (identical or near-duplicate reformulations). Table 4 summarizes trajectory properties, and Table 5 reports their usage. We interpret trajectories with three stability measures: Dense Similarity is cosine similarity between consecutive query embeddings (Qwen3-Embedding-0.6B (Zhang et al., 2025)); Jaccard Similarity is lexical overlap over lowercase, whitespace-tokenized word sets; and Result Overlap is the Jaccard overlap between retrieved document identifier sets for consecutive queries (definitions in Appendix B).
The “Drill-Down” Bias:
Across intents, agents mainly tighten constraints via local edits or pivot across nearby facets, while explicit broadening is consistently the least-used move (under 11%; Table 5). This imbalance suggests that agents are more comfortable focusing on a local neighborhood of the query space rather than stepping back to relax constraints and reconsider alternatives. Exploration is common (roughly 36–48%), but pivots tend to induce larger evidence turnover and slower transitions, making them costlier than incremental refinement (Table 4). Together, these signatures suggest that when progress stalls, agents often continue local edits rather than deliberately broadening and re-planning, motivating improved controller policies and training signals.
Intent Differences:
Although all intents share the same move vocabulary, they exhibit different interaction patterns (Table 5). Declarative sessions are most prone to retry-like behavior, with Repetition at about one-third of moves, consistent with agents re-issuing near-duplicate queries when evidence remains elusive. Reasoning sessions, in contrast, sustain the highest pivoting (Exploration near 48%) with lower retrying, suggesting a broader search over sub-questions. Procedural sessions more often combine pivots with subsequent constraint tightening, aligning with an “explore then refine” workflow in step-by-step tasks.
Figure 4. Step-wise trajectory distribution trends for the first 10 steps across different task intents. Each sub-figure illustrates the evolving proportions of Specialization, Generalization, Exploration, and Repetition as the search session progresses.
Figure 5. A Declarative retry-loop example dominated by near-duplicate reformulations.
Stability as a Diagnostic:
[…]
Pacing Implications:
Move types also differ in end-to-end pacing between requests (Table 4). Exploration is slower (median of 14.0s) than minor reformulations such as Repetition (6.0s), consistent with facet pivots inducing larger evidence turnover and higher processing cost. This makes strategy selection consequential, when pivots are expensive, agents may default to cheaper local edits unless they learn when refinement is no longer productive. Although explicit broadening is rare, it is often followed by re-specialization in the transition dynamics (Section 7.2), consistent with broadening acting as a brief reset rather than sustained re-planning.
Figure 6. Trajectory transition matrix (row-wise normalized).
7.2. Temporal Dynamics of Search Strategies
We next study how query-reformulation strategies evolve over a session. Figure 4 traces the step-wise trajectory composition over the first 10 steps for each intent, and Figure 6 summarizes how moves transition from one step to the next.
Trends over Steps:
Strategy use shifts over time. Early steps mix facet pivots, retries, and constraint adjustments, then diverge by task type (Figure 4). Declarative sessions gradually concentrate on retries, consistent with late-stage stagnation; Procedural sessions maintain substantial pivoting but increasingly emphasize refinement; and Reasoning sessions sustain pivoting with consistently low retrying. We report the full step-wise proportions in Figure 4 and focus here on higher-level directional shifts.
Transition Mechanisms:
Figure 6 helps explain these trends by showing which moves persist as runs and which act as resets. Exploration and Repetition often form multi-step runs, consistent with the step-wise patterns in Figure 4 where pivoting and retrying persist across consecutive steps. In contrast, broadening frequently acts as a brief reset, nearly half of Generalization moves are followed by Specialization, suggesting that agents relax constraints momentarily before re-introducing them.
Case Study - The “Reset-then-Refine” Pattern:
Figure 7 illustrates a reset-then-refine sequence. The agent specializes a broad topic by adding constraints (Napoleon campaigns $\rightarrow$ Italy 1796), then generalizes by removing them (a shorter, broader query), and re-specializes toward a different facet (Egyptian expedition). Query-length changes match our definitions, where Specialization tends to lengthen queries, while Generalization shortens them (Table 4). This is consistent with Generalization acting as lightweight backtracking to switch refinement branches rather than sustained broadening.
Figure 7. A reset-then-refine example: Specialization $\rightarrow$ Generalization $\rightarrow$ Specialization.
7.3. Context-Driven Term Adoption Rate (CTAR)
The previous sections show that agents frequently pivot, refine, and retry across steps, and that these strategies shift over a session’s lifespan. A central question is whether such multi-step behavior is evidence-grounded: if agents do not incorporate retrieved context when reformulating queries, multi-step search may degenerate into a sequence of effectively independent single-shot queries. Thus, we measure context traceability by testing whether new query terms introduced at step k+1k{+}1 appear in the evidence context up to step $k$. We use CTAR as defined in Section. 4.2 and report both last-step and aggregated variants.
Evidence of Context Grounding and Recency Effect:
The CTAR analysis indicates that a substantial fraction of newly introduced query terms can be lexically traced to retrieved evidence. Table 6 reports the mean CTAR under aggregated versus last-step context.
Table 6. Mean CTAR under Aggregated vs. Last-step Evidence.
Trajectory type Aggregated Last-step
Specialization 78.35% 70.93%
Generalization 52.95% 49.44%
Exploration 69.59% 60.14%
Repetition 20.92% 19.77%
Overall 54.35% 48.54%
Key observations are that overall more than half of newly introduced query terms are present in the aggregated evidence context (mean CTAR = 54.35%), aggregated context adds +5.81 percentage points over last-step evidence suggesting strong reliance on recent evidence with additional benefit from earlier steps (Liu et al., 2024), and CTAR varies substantially by trajectory type, with Specialization and Exploration much higher than Repetition (78.35%/69.59% vs. 20.92% under aggregated context).
CTAR is intentionally a lexical traceability measure (based on token overlap), rather than a semantic variant, to avoid embedding-model dependence and uninterpretable similarity-threshold choices. As a result, a low CTAR score does not necessarily imply that the agent ignores evidence. Conversely, a high CTAR score indicates that new terms are explicitly present in the context, but it does not by itself establish causality.
Multi-step Context Contribution:
We further examine how CTAR varies when tracing new terms against evidence from different historical steps within a session. Figure 8 summarizes the CTAR measured against progressively older evidence contexts. For example, a value of 70.93% for Spec at “previous step 1” means that, for Specialization transitions, 70.93% of newly introduced terms in $q_{k+1}$ also appear in the evidence retrieved at the immediately preceding step $E_{k}$ (i.e., the Last Step context in Table 6); “previous step 2” analogously traces against $E_{k-1}$, and so on. The curves suggest that for trajectory types with higher CTAR, term adoption remains non-trivial beyond the immediate previous step, indicating that historical context can contribute to query reformulation, in addition to the most recent evidence.
Figure 8. CTAR scores across previous steps.
Taken together, these results suggest that the agent’s query reformulation behavior is frequently consistent with evidence-grounded term adoption, where new query terms are often directly traceable to retrieved context, with a strong reliance on the most recent step and additional contributions from earlier steps.
8. Discussion and Implications for Agent Design
Our findings suggest that agentic search is not merely a sequence of queries but a structured process of state transitions and evidence integration, with three implications for agent design:
Repetition as a Potential Stall Signal:
In Declarative sessions, repetition increases to 42.68% by Step 9 (Figure 4), indicating that agents can enter near-duplicate loops when retrieved evidence does not diversify. Repetition can therefore be treated not only as redundancy but also as a practical stall signal. Systems may benefit from early-detection guards that trigger a strategy switch (e.g., forcing a Generalization move) or escalation to a human-in-the-loop when lexical repetition exceeds a threshold (Feild et al., 2010; Teevan et al., 2007).
Intent-Adaptive Resource Allocation:
Retrieval depth is largely rigid, with 91.64% of requests using $K\in\{1,5,10\}$, despite intent-dependent needs (e.g., deeper retrieval for Procedural than Declarative). This suggests retrieval is often treated as a static hyperparameter. Future architectures may adopt intent-aware budgeting that adjusts compute and retrieval depth across intents and steps, rather than relying on fixed $K$ choices (Jeong et al., 2024).
Evidence Grounding as an Audit Signal:
The gain in CTAR from aggregated context (+5.81 pp over the last-step) shows that agents actively synthesize historical evidence. Crucially, the sharp contrast in grounding between progress-oriented moves like Specialization (78.35% CTAR), and moves with lower CTAR like Repetition (20.92% CTAR), suggests that low evidence adoption is associated with retry loops. This motivates explicit context management modules that not only cache history but prioritize high-utility terms from earlier steps, to guide subsequent query formulation and reduce ungrounded exploration (Trivedi et al., 2023).
9. Conclusions
We study agentic search behavior in the wild through 14.44M DRGym requests (Coelho et al., 2025), converting raw logs into sessions and analyzing session-level structure and step-wise query transitions. Our central takeaway is that multi-step agentic search exhibits intent-conditioned reformulation patterns that expose where agents make progress versus stall (Sections 57). Importantly, our diversity analyses in Section 3 show that queries from this log is not simply repetitions of popular benchmarks; instead, they cover a broad range of semantic content, supporting the ecological validity of the behavioral patterns we report. In particular, trajectories reveal a drill-down bias. Agents favor local refinement and facet pivots over deliberate broadening or backtracking. Declarative sessions exhibit salient retry loops, while Procedural and Reasoning sessions show distinct progression signatures (Sections 67). To study how multi-step querying incorporates retrieved content without clicks, we introduce CTAR and find that newly introduced query terms often align with retrieved context across steps, suggesting non-trivial cross-step evidence reuse (Section 7.3).
These results suggest several implications for designing more reliable agentic IR systems. First, the move distributions, transition dynamics, and case studies in Section 7 highlight a recurring pattern where agents often overuse inexpensive local edits rather than switching strategies. This motivates controller policies that detect high-stability loops and trigger structured interventions (e.g., forced facet pivots, purposeful backtracking, or adaptive retrieval budgeting). Second, our intent-conditioned analyses indicate that a single global reformulation policy may be suboptimal; intent-aware guardrails can better balance refinement, exploration, and retrying under different tasks (Section 6). Third, CTAR offers a lightweight audit signal for whether multi-step querying is evidence-driven, and can support memory designs that preserve useful context while reducing redundant retries (Section 7.3).
Finally, we see several promising directions for follow-up work enabled by this study. Beyond releasing the DRGym-derived dataset and the analysis protocol, future research can connect reformulation moves to downstream answer quality, develop learning objectives that explicitly reward productive backtracking and evidence-grounded progress, and explore how retrieval parameters and context management should adapt over the course of a session. More broadly, we hope these measurements and case-driven diagnostics provide a foundation for intent-aware control and training of agentic search systems under reproducible retrieval settings.
Appendix A Log Sessionization Procedure
This section describes our log sessionization pipeline, including the semantic continuity model used to link adjacent queries and the per-IP online assignment rules used to form sessions.
Step 1: Train a semantic continuity model.
(1) Training pairs:
We randomly sample $\sim$200K queries and pair each query with the nearest-in-time query from the same IP.
(2) Pair labels: We label each pair as same-session vs. different-session using an LLM-as-a-judge prompt with gpt-5-nano-2025-08-07 (Appendix D.1).
(3) Pair representation: We encode each query with Qwen3-Embedding-0.6B (Zhang et al., 2025) and use the resulting embeddings to construct a fixed dense feature vector for each query pair as input to the downstream neural classifier.
(4) Classifier: We train a 3-layer MLP to output a continuity score in $[0,1]$. The hidden-layer dimensions are 1024, 512, and 256, and the model achieves a held-out accuracy of 0.9419.
Step 2: Sessionize online per IP (with validation).
(1) Per-IP assignment:
We process queries in order and maintain active sessions for each IP. For an incoming query $q_{t}$, we score it against each active session using the session’s most recent query and assign $q_{t}$ to the highest-scoring session if the score $\geq 0.5$; otherwise we start a new session.
(2) Temporal hard cutoff: Beside the continuity score, if the gap to the candidate session’s last query exceeds 10 minutes, we start a new session.
(3) Sanity check: We manually inspect 100 random sessions for coherence; after excluding four unusually long sessions, the remaining sessions are centered on a single objective.
Appendix B Auxiliary Metric Definitions
This section defines the auxiliary metrics used throughout our analyses and provides their formal notation and formulas for reproducibility (Table 7).
Notation:
For a session $s=(q_{1},\ldots,q_{|s|})$, let $\mathbf{v}_{t}$ denote the dense embedding of query $q_{t}$ (and $\cos(\cdot,\cdot)$ the cosine similarity). $W_{t}$ is the set of normalized tokens from $q_{t}$ after lowercasing and stopword-aware tokenization $\mathrm{WS\_tok}(\cdot)$. $D_{t}$ is the set of retrieved evidence returned for query $q_{t}$ (at the logged retrieval depth).
Table 7. Summary of auxiliary metrics used in our analyses.
Metric Formula
Initial–Final Gap $\mathrm{Gap}(s)=1-\cos\!\big(\mathbf{v}_{1},\ \mathbf{v}_{|s|}\big)$
Dense Sim. $\mathrm{DenseSim}(q_{t},q_{t+1})=\cos\!\big(\mathbf{v}_{t},\mathbf{v}_{t+1}\big)$
Jaccard Sim. $\mathrm{Jac}(q_{t},q_{t+1})=\frac{|W_{t}\cap W_{t+1}|}{|W_{t}\cup W_{t+1}|},\qquad W_{t}=\mathrm{WS\_tok}(\mathrm{lower}(q_{t}))$
Result Overlap $\mathrm{Overlap}(q_{t},q_{t+1})=\frac{|D_{t}\cap D_{t+1}|}{|D_{t}\cup D_{t+1}|}$
Appendix C Representative Query Examples
This section presents real representative query examples from our logs to help interpret our intent (Table 8) and trajectory (Table 9) labels.
Table 8. Representative Queries for Intent Categories.
Intent Example Queries
1. Who owns Handi-Snacks?
Declarative 2. Definition of home food store
3. Ansel Adams residences in Yosemite National Park
1. Best instructions for homemade reading shelf
Procedural 2. Practical driving tips for beginners
3. How to change a constitution?
1. Why is gas so expensive?
Reasoning 2. Does advertising help or harm us?
3. Why are minority rights important?
Table 9. Representative Transitions Trajectories. Each entry illustrates a step-wise reformulation ($q_{k}\rightarrow q_{k+1}$).
Type Step Example Queries
Specialization 1 Recent climate data Durban
2 Average temperature in Durban 2025
Generalization 1 Key events that ended Hitler’s dictatorship
2 Hitler’s dictatorship
Exploration 1 Headquarters of Oberoi Hotels
2 Parent company of Oberoi Hotels
Repetition 1 Handi-Snacks parent company
2 Who owns Handi-Snacks
Appendix D LLM-as-a-judge Prompts and Parsing Details
This section provides the exact LLM-as-a-judge prompts for reproducibility.
D.1. Query-pair Continuity Judgment Prompt
[SYSTEM]
You label query pairs for a DeepResearch search agent. The agent fans out several queries to answer ONE user question.
Answer YES if both queries would naturally be used for the same research task or user question (same core topic), even if they cover different aspects or levels of detail.
Answer NO if they would answer clearly different questions, even if they share broad words like WWII’, health’, or economy’.
[USER]
Query 1: <<query1>>
Query 2: <<query2>>
For a DeepResearch agent that fans out queries to answer ONE user question,
would these two queries belong to the same research task?
Answer YES or NO only.
D.2. Session-level Intent Classification Prompt
[SYSTEM]
You are an expert search intent classifier.
[USER]
"Session Queries:\
<<joined_queries>>
Classify the user intent of this session into exactly ONE of these three categories:
1. Declarative: Asking for simple facts, definitions, entity attributes, or lists (e.g., who is’, what is’, release date’).
2. Procedural: Asking for steps, methods, tutorials, or guides (e.g., how to’, guide for’, fix error’).
3. Reasoning: Asking for comparisons, planning, analysis, multi-hop reasoning, or creative generation (e.g., difference between’, best plan for’, why is’).
Output ONLY the category name (Declarative, Procedural, or Reasoning).
D.3. Step-wise Trajectory Classification Prompt
[SYSTEM]
You are an expert search behavior analyst.
[USER]
Query 1 (Previous): <<PLACEHOLDER: q_k>>
[…]
Analyze the search behavior evolution from Query 1 to Query 2 for an autonomous agent.
Classify the transition into exactly ONE of these four categories:
1. Specialization (Vertical Deepening): Query 2 is MORE specific than Query 1 by adding constraints/details (q2 $\subset$ q1). (e.g., apple -> green apple nutritional value’).
2. Generalization (Vertical Broadening): Query 2 is MORE general than Query 1 by removing constraints/abstracting (q2 $\supset$ q1). (e.g., green apple nutritional value -> benefits of fruits’).
3. Exploration (Horizontal Expansion within the same domain/task): Query 2 is NOT simply more specific or more general. It shifts to a different aspect/subtopic/related entity but still within the same overall topic/domain. (e.g., green apple nutritional value -> green apple recipes or MRI Scans -> CT Scans’).
4. Repetition (Stationary): Query 2 is semantically equivalent to Query 1. It is a paraphrase, reformatting, or synonym replacement with NO significant change in intent (e.g., green apple value -> nutritional value of green apple’).
Output ONLY the category name (Specialization, Generalization, Exploration, or Repetition).